OpenStudy (anonymous):

agnes has a weighted coin so that heads come up 2/3 of the time.she tosses her coin five times. what is the probability that heads will appear for the first time on toss 5

OpenStudy (agent0smith):

If heads appears for the first time on toss 5, then the first four tosses must all be tails. The probability of tails is 1-(2/3) = 1/3.

OpenStudy (agent0smith):

So the first four tosses must all be tails, then toss 5 is heads: \[P = \frac{ 1 }{ 3 } \times \frac{ 1 }{ 3 } \times \frac{ 1 }{ 3 } \times \frac{ 1 }{ 3 } \times \frac{ 2 }{ 3 } \]

OpenStudy (anonymous):

that what i did

OpenStudy (anonymous):

q^1-i * p which p=2/3 q=1-2/3 i=5 then (1/3)^4 * 2/3